Lin Corporation has a single product whose selling price is $134 per unit and whose variable expense is $67 per unit. The company’s monthly fixed expense is $32,300. Required: 1. Calculate the unit sales needed to attain a target profit of $7,900. (Do not round intermediate calculations.) 2. Calculate the dollar sales needed to attain a target profit of $8,600. (Round your intermediate calculations to the nearest whole number.)
Answer #1

Contribution margin=Sales-Variable cost  

=(134-67)=$67 per unit

Contribution margin ratio=Contribution margin/Sales

=(67/134)=0.5

a.Target Contribution margin=Fixed expenses+Target profit

=(32300+7900)=$40200

Hence target sales=(Target Contribution margin/Contribution margin per unit)

=40200/67

=600 units

b.Target Contribution margin=Fixed expenses+Target profit

=(32300+8600)=$40900

Hence target sales=(Target Contribution margin/Contribution margin ratio

=40900/0.5

=$81800